My hotes Asa Your consider the approximately normal population of heights of male college students with mean-68 inches and standard deviation of σ-43 inches. A random u ple obtained igts (a) Describe the distribution of x, height of male college students skewed right approximately normal skewed lef chi-square b) Find the proportion of male college students whose height is greater than 69 inches. (Give yeur answer correct to four decimal places) (c) Describe the distribution of x, the mean of samples of size 12 O skewed right approximately normal skewed left O chi-square (d) Find the mean of the distribution. (Give your answer correct to the nearest whole number.) Find the standard error of the distribution. (Give your answer correct to two decimal places) (e) Find P>71) (Give your answer correct to four decimal places.)
Answer #1

a)

approximately normal

b)

P(X>69)=P(Z>(69-68)/4.3)=P(Z>0.23)=0.4090

c)

approximately normal

d)mean =68

ii)std error =4.3/sqrt(12)=1.24

e)P(Xbar>71)=P(Z>(71-68)/1.24)=P(Z>2.42)=0.0078